OpenStudy (anonymous):

How can I find the values of x and y in an isosceles triangle in the drawing I will show...please help!!!

OpenStudy (anonymous):

 Please help!

OpenStudy (anonymous):

4y is the vertex angle and 2x + y and 79- x are the base angles

OpenStudy (anonymous):

Set 2x + y = 79- x

OpenStudy (anonymous):

@middi1997 Helloo..o!

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

2x + y = 79- x

OpenStudy (anonymous):

Gather x and y terms to the left side, will you!

OpenStudy (anonymous):

ok h/o

OpenStudy (anonymous):

I got 3x + y = 79

OpenStudy (anonymous):

Then the second equation is total of angles in triangle = .... degree?

OpenStudy (anonymous):

um I do not understand

OpenStudy (anonymous):

oh sorry its 180 degrees

OpenStudy (anonymous):

Yup :) Now add all these 3 angles together, can you?

OpenStudy (anonymous):

well what do i do about 79 would it be 3x + 5y = 180

OpenStudy (anonymous):

( 2x +y) + ( 79 -x ) + 4y = ...?

OpenStudy (anonymous):

I mean it's not 3x!

OpenStudy (anonymous):

= 180 but what do we do with the 3x + y

OpenStudy (anonymous):

wait im confused sorry

OpenStudy (anonymous):

Focus on the left side first!

OpenStudy (anonymous):

so it starts of like this: 2x + y = 79 - x 2x + y + x = 79 - x + x 3x + y + 79 and now im confused

OpenStudy (anonymous):

This is the first equation, correct!

OpenStudy (anonymous):

But we need the second one to solve for both x, y!

OpenStudy (anonymous):

so the second one is 2x +y) + ( 79 -x ) + 4y? but why

OpenStudy (anonymous):

So we make the second one from total of angles = 180!

OpenStudy (anonymous):

As a thumb rule, when we have 2 unknowns, we need 2 equations !

OpenStudy (anonymous):

ok so the result of the second equation i have so far is x + 5y = 101

OpenStudy (anonymous):

Correct now :) 3x + y = 79 (1) x + 5y = 101 (2) Can you solve system of 2 equations to find x, y?

OpenStudy (anonymous):

I don't think I am that advanced yet

OpenStudy (anonymous):

gaah im in geometry but i never had algebra 1 or 2

OpenStudy (anonymous):

oh ok do you know a short way to explain system of equations

OpenStudy (anonymous):

However, don't be discourage, I'll show you if you want to learn :)

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

ok so -3(3x + y) = -3(79) y = - 237

OpenStudy (anonymous):

Distribute -3 in: -3 ( x + 5y = 101) -3x + .. = ....

OpenStudy (anonymous):

oh to make them equal so we can subtract ( well not necessarily =) so it is 3x + y = 79 -3x - 15y = -303

OpenStudy (anonymous):

-14y = -224 ------------ -14

OpenStudy (anonymous):

=16 y = 16

OpenStudy (anonymous):

so now can we substitute y in 3x + y = 79 to be 3x + 16 = 79

OpenStudy (anonymous):

3x + 16 = 79 - 16 -16 3x= 63 --------- 3 x = 21

OpenStudy (anonymous):

thank you so much... but i have a second problem i need he;p with if you are willing

OpenStudy (anonymous):

Typo x = 21 degree!

OpenStudy (anonymous):

Not yet, the last step is replace x, y values into the angles: -> 4y = 64 degree ....

OpenStudy (anonymous):

oh no i wasnt solving for degrees just for the values of x and y

OpenStudy (anonymous):

but will you help me with my next problem?

OpenStudy (anonymous):

Oh, so it's done :) To make it simple, open new post per question :)

OpenStudy (anonymous):

ok i will h/o

OpenStudy (anonymous):

Make sure you understand the concept prior to moving on the next level!